OpenStudy (anonymous):

Glenda began the day with a golf score of -6 and ended with a score of -10. Which statement represents her golf score for that day?

OpenStudy (anonymous):

Glenda's original score decreased by -4

OpenStudy (anonymous):

A. -6 - (-10) =4 B. -10 - (-6)= -4 C. -6 + (-10)= -16 D. -10 + (-6) = -16

OpenStudy (anonymous):

since her score decreased by 4 points it is b.-10-(-6)= -4